OpenStudy (anonymous):

A bicycle wheel with radius 26" rotates through an arc that measures 80°. What is the length of the arc of the tire that touched the ground? a. 2.88 pi inches b. 5.78 pi inches c. 11.56 pi inches

OpenStudy (anonymous):

Arc = (degree/360)* circumference

OpenStudy (nurali):

Arc Length=\[\frac{ x }{ 360 }*2pir\]

OpenStudy (nurali):

If x = 80, and r = 26, then Arc length = (x/360) . 2pi.r = (80/360).2pi(26) = ....

OpenStudy (anonymous):

I got 5.78 pi